# Multimodal-Wireless: A Large-Scale Dataset for Sensing and Communication

### v1

@@ Abstract
This paper presents Multimodal-Wireless, an open-source multimodal sensing dataset designed for wireless communication research. The dataset is generated through an integrated and customizable data pipeline built upon the CARLA simulator and Sionna framework. It contains approximately 160,000 frames collected across four virtual towns, sixteen communication scenarios, and three weather conditions, encompassing multiple sensing modalities—communication channel, light detection and ranging, RGB and depth cameras, inertial measurement unit, and radar. This paper provides a comprehensive overview of the dataset, outlining its key features, overall framework, and technical implementation details. In addition, it explores potential research applications concerning communication and collaborative perception, exemplified by beam prediction using a multimodal large language model. The dataset is open in https://le-liang.github.io/mmw/.
@@ I Introduction
While numerous datasets have been developed for intelligent autonomous systems [4, 5, 6, 7, 8, 9, 10], a critical dichotomy exists. Datasets for collaborative perception (e.g., OPV2V [4], DAIR-V2X [5]) excel in sensory modalities like cameras and light detection and ranging (LiDAR), whereas wireless communication datasets (e.g., BostonTwin [6], DeepMIMO [7]) focus on channel modeling. The integration of these two domains remains rare. Pioneering efforts to combine them, such as E-Flash [8], ViWi [9], and DeepSense 6G [10], lack crucial elements like adverse weather conditions, high-frequency channel state information (CSI), or complete multimodal synchronization. As summarized in Table I, these gaps reveal a clear need for a unified, weather-resilient, and extensible dataset to advance robust multimodal communication research.
@@ IV Multimodal-Wireless: Detailed Specifications
The integration of the CARLA simulator, Blender and Sionna facilitates the synchronization between traditional modalities and communication channel modality, as well as the spatial consistency of the CARLA world and Sionna scenes. To realize the pipeline, we need to let the scenes flow smoothly among these three platforms. In this section, we clarify the parametric and technical details in the CARLA simulator, Blender and Sionna, respectively.
@@ IV-A CARLA–Scenario Execution and Data Collection
LiDAR
64 channels, $30k$ points per sample,
$120$ m capturing range, $25^{\circ}$ to $2^{\circ}$
vertical FOV
IMU
Gyroscope noise: mean $0.001$ rad/s,
standard deviation (std) $0.002$ rad/s,
accelaration noise: std $0.1$ m/s2
Radar
$2k$ points per sample, $100$ m capturing
range, $30^{\circ}$ vertical FOV and $110^{\circ}$
horizontal FOV
@@ IV-B Blender–Scenario Reconstruction
Ground Sunny Weather
itu_very_dry_ground
itu_medium_dry_ground
@@ IV-C Sionna–Channel Generation
where $\mathbf{H}(f_{k})$ is the channel frequency response at the $k$-th subcarrier. Here, $f_{k}$ is the baseband frequency of the $k$-th subcarrier and satisfies $f_{k}=(k-\frac{K+1}{2})\Delta f$ for an OFDM system with $K$ subcarriers and $\Delta f$ subcarrier spacing. All simulation parameters of Multimodal-Wireless are summarized in Table V. We categorize them into two groups: ray-tracing parameters, which are set during the physics-based ray-tracing process, and communication parameters, which can be customized by the user when synthesizing the frequency-domain channel from the raw path data ($\{\mathbf{A}_{m},\tau_{m}\}_{m=1}^{M}$). This mechanism grants users maximum control over the final channel realization.
@@
Ray-tracing
Carrier Frequency $28$ GHz / $4.9$ GHz
@@
Communication
Subcarrier Spacing $120$ kHz
@@ V Example Research Applications
As a case study, we present multimodal large language model (LLM) based beam prediction. Specifically, we predict future beam indices of $W=10$ time steps (i.e., $100$ ms) with history beam indices, LiDAR and RGB camera data of $P=40$ time steps (i.e., $400$ ms). We utilize the data collected under the “sunny” weather, with a training set of 43,040 samples and a validation set of 5,380 samples constructed. We employ a pre-trained GPT-2 model as the LLM backbone. For modality-alignment, the 10 Hz RGB and LiDAR data are zero-padded to synchronize with the 100 Hz beam index sequence. BS and user apply $1\times 16$ ULAs. We select the combiner $\mathbf{w}$ and the precoder $\mathbf{f}$ from a $Q$-DFT codebook, $Q$ set to 64, where the $q$-th candidate beamformer is expressed as
@@
As shown in Fig. 4, by incorporating LiDAR and camera data, our multimodal LLM-based method consistently achieves higher normalized gain than the method in [16] across all prediction time steps, addressing the absence of perfect AOD on the BS side in practice. The signal power after beamforming increases by an average of 0.15 dB for the final five time steps. This improvement arises because historical LiDAR and camera data provide valuable environmental context, allowing the model to better capture scene dynamics such as incoming traffic jams. Fig. 5 reveals that combining LiDAR with beam index data achieves performance close to the full-modality model. In contrast, retaining only the camera input leads to a degradation, which is owing to the limited FOV of the RGB camera that compromises the quality of the camera modality data. However, compared with the index-only baseline, the addition of the camera modality still improves performance.

### v2

@@ Abstract
This paper presents Multimodal-Wireless, a large-scale open-source dataset for multimodal sensing and communication research. The dataset is generated through an integrated and customizable data pipeline built upon the CARLA simulator and Sionna framework, and features high-resolution communication channel state information (CSI) fully synchronized with five other sensor modalities, namely LiDAR, RGB and depth camera, inertial measurement unit (IMU) and radar, all sampled at 100 Hz. It contains approximately 160,000 frames collected across four virtual towns, sixteen communication scenarios, and three weather conditions. This paper provides a comprehensive overview of the dataset, outlining its key features, overall framework, and technical implementation details. In addition, it explores potential research applications concerning communication and collaborative perception, exemplified by beam prediction using a multimodal large language model. The dataset is open in https://le-liang.github.io/mmw/.
@@ I Introduction
While numerous datasets have been developed for intelligent autonomous systems [4, 5, 6, 7, 8, 9, 10], a critical dichotomy exists. Datasets for collaborative perception (e.g., OPV2V [4], DAIR-V2X [5]) excel in sensory modalities like cameras and LiDAR, whereas wireless communication datasets (e.g., BostonTwin [6], DeepMIMO [7]) focus on channel modeling. The integration of these two domains remains rare. Pioneering efforts to combine them, such as E-Flash [8], ViWi [9], and DeepSense 6G [10], lack crucial elements like adverse weather conditions, high-frequency channel state information (CSI), or complete multimodal synchronization. As summarized in Table I, these gaps reveal a clear need for a unified, weather-resilient, and extensible dataset for multimodal communication research.
@@ IV Multimodal-Wireless: Detailed Specifications
The integration of CARLA, Blender and Sionna facilitates the synchronization between traditional modalities and communication channel modality, as well as the spatial consistency of the CARLA world and Sionna scenes. To realize the pipeline, we need to let the scenes flow smoothly among these three platforms. In this section, we clarify the parametric and technical details in CARLA, Blender and Sionna, respectively.
@@ IV-A CARLA–Scenario Execution and Data Collection
LiDAR 64 channels, $30k$ points per sample, $120$ m capturing range, $25^{\circ}$ to $2^{\circ}$ vertical FOV
IMU Gyroscope noise: mean $0.001$ rad/s, standard deviation (std) $0.002$ rad/s, accelaration noise: std $0.1$ m/s2
Radar $2k$ points per sample, $100$ m capturing range, $30^{\circ}$ vertical FOV and $110^{\circ}$ horizontal FOV
@@ IV-B Blender–Scenario Reconstruction
Ground Sunny Weather itu_very_dry_ground itu_medium_dry_ground
@@ IV-C Sionna–Channel Generation
where $\mathbf{H}(f_{k})$ is the channel frequency response at the $k$-th subcarrier. Here, $f_{k}$ is the baseband frequency of the $k$-th subcarrier and satisfies $f_{k}=(k-\frac{K+1}{2})\Delta f$ for an OFDM system with $K$ subcarriers and $\Delta f$ subcarrier spacing. Simulation parameters of the dataset are summarized in Table V, and categorized into two groups: ray-tracing parameters, which are set during the physics-based ray-tracing process, and communication parameters, which can be customized by the user when synthesizing the frequency-domain channel from the raw path data ($\{\mathbf{A}_{m},\tau_{m}\}_{m=1}^{M}$). This mechanism grants users maximum control over the final channel realization.
@@
Ray-tracing Carrier Frequency $28$ GHz / $4.9$ GHz
@@
Communication Subcarrier Spacing $120$ kHz
@@ V Example Research Applications
As a case study, we present multimodal large language model (LLM) based beam prediction. Specifically, we predict future beam indices of $W=10$ time steps (i.e., $100$ ms) with history beam indices, LiDAR and RGB camera data of $P=40$ time steps (i.e., $400$ ms). We utilize the data collected under the “sunny” weather, with a training set of 43,040 samples and a validation set of 5,380 samples constructed. We employ a pre-trained GPT-2 model as the LLM backbone. For modality-alignment, the 10 Hz RGB and LiDAR data are replicated to synchronize with the 100 Hz beam index sequence. BS applies a ULA of size $1\times 64$ and user of size $1\times 16$. We select the combiner $\mathbf{w}$ and the precoder $\mathbf{f}$ from a $Q$-DFT codebook, $Q$ set to 64, where the $q$-th candidate beamformer is
@@
As shown in Fig. 4, by incorporating LiDAR and camera data, our multimodal LLM-based method achieves higher normalized gain than the method in [16] across all prediction time steps, addressing the absence of AOD on the BS side in practice. This improvement arises because historical multimodal data provide valuable environmental context, allowing the model to better capture scene dynamics such as incoming traffic jams. Fig. 5 reveals that combining LiDAR with beam index data achieves performance close to the full-modality model. In contrast, retaining only the camera input leads to a degradation, which is owing to the limited FOV of the RGB camera that compromises the quality of the camera modality data. However, compared with the index-only baseline, the addition of the camera modality still improves performance.
@@ VI Conclusion
Acknowledgement
The work was supported by the National Key R&D Program of China under Grant 2024YFE0200700, and the National Natural Science Foundation of China (NSFC) under Grant W2421087.